need help solving this problem:
A=18.98 B=3.3 C=2.28 solve for D
A + B x C >= D
thanks.

Answer by arunpaul(104)

A=18.98, B=3.3, C=2.28
A + B x C >= D
substituting the value of A,B,C in the above equation we have
18.98+3.3 X 2.28 >= D
as per the rule of equation
multiplication operation will be done first so the above equation will be
18.98+(3.3 X 2.28) >= D
18.98 + 9.24 >= D
28.22 >= D
so D must be <= 28.22 (Ans)